OpenStudy (anonymous):

What is the approximate distance from (-2, 9) to (1, 4)? Answer 5 units 8 units 6 units 7 units (I chose 7 units is that right?)

OpenStudy (katrinakaif):

Can you place these on a coordinate plane for me first?

OpenStudy (anonymous):

there is no coordinate plane sorry

OpenStudy (katrinakaif):

Do you know the distance formula?

OpenStudy (anonymous):

\[\sqrt{3^2+5^2}=\sqrt{9+25}=\sqrt{34}\] closes whole number is 6